OpenStudy (anonymous):

how do i simplify this : x*x^(1/3)*x^(-2/3)

OpenStudy (kc_kennylau):

You mean simplify? Please don't abuse the verb "solve".

OpenStudy (anonymous):

Yes, sorry! simplify

OpenStudy (kc_kennylau):

a^n*a^m = a^(n+m)

OpenStudy (anonymous):

x*x^(1/3+(-2/3), right?

OpenStudy (kc_kennylau):

Yep :)

OpenStudy (kc_kennylau):

Don't forget that x = x^1 :)

OpenStudy (anonymous):

ahhh , so it's x^(1+(1/3)+(-2/3)

OpenStudy (kc_kennylau):

and simplify the mess inside :)

OpenStudy (anonymous):

Great! Thanks! :D

OpenStudy (kc_kennylau):

No problem :)